# New Primitive OBZCPs of Lengths up to 49

### v1

@@ Abstract
This paper aims to search for new primitive optimal and sub-optimal Odd Binary Z-Complimentary Pairs (OBZCPs) for lengths up to 49. As an alternative to the celebrated binary Golay complementary pairs, optimal OBZCPs are the best almost-complementary sequence pairs having odd lengths. We introduce a computer search algorithm with complexity $O(2^{N})$, where $N$ denotes the sequence length and then show optimal results for all $27\leq N\leq 33$ and $N=37,41,49$. For those sequence lengths (i.e., $N=35,39,43,45,47$) with no optimal pairs, we show OBZCPs with largest zero-correlation zone (ZCZ) widths (i.e., $Z$-optimal). Finally, based on the Pursley-Sarwate criterion, we present a table of OBZCPs with smallest demerit factors in terms of the combined auto-correlation and cross-correlation.
@@ I-A Background
Complementary pairs of sequences are a useful concept in coding theory, wireless communication, radar sensing, and signal processing. The general design objective is to find two equal-length sequences whose maximum aperiodic auto-correlation function (AACF) sums should be as small as possible (ideally zero). Pioneered by Marcel J. E. Golay [1] pioneered in 1951, binary complementary pair was first studied in his design of infrared multislit spectrometry, a detector which isolates the desired radiation with a fixed single wavelength from background radiation with many different wavelengths. Formally, a pair of sequences is called a Golay complementary pair (GCP) [2] if their AACF sums are zero for all the non-zero time-shifts. Since it is generally difficult to find a single unimodular sequence with zero AACF sidelobes [^1] Although a Huffman sequence has zero AACF sidelobes except for the end time-shift, its sequence elements may have different magnitudes., GCP provides a solution by allowing two sequences to work in a collaborative way. Some representative applications of the GCPs (and their extensions/variants) include: peak-to-mean envelope power ratio (PMEPR) reduction of multicarrier signals [3, 4, 5, 6], Doppler resilient radar waveforms [7, 8, 9], channel estimation [10, 11, 12, 13], intercell interference rejection [14], multicarrier code-division multiple access [15, 16, 17], etc.
Despite their wide applications, however, binary GCPs are limited to even-lengths only. More specifically, the existing binary GCPs are only known to have sequence lengths of $2^{\alpha}10^{\beta}26^{\gamma}$ (where $\alpha,\beta,\gamma$ are non-negative integers). For the odd-length case, optimal binary almost-complementary pairs are studied in [18]. Their optimization criteria is to maximize the zero-correlation zone (ZCZ) width of an odd-length binary pair, whilst minimizing its out-of-zone AACF sums. Such optimal binary pairs exhibit the correlation properties closest to GCPs and are called optimal odd-length binary Z-complementary pairs (OBZCPs). It is found that each optimal OBZCP has maximum ZCZ width of $(N+1)/2$, and minimum out-of-zone AACF sum magnitude of 2, where $N$ denotes the sequence length (odd).
@@ I-B Contributions
Searching for new optimal OBZCPs is of interest for understanding their deeper structural properties and for providing higher level of flexibility in practical applications. Besides the aforementioned systematic constructions, it is equally important to look for short and primitive OBZCPs because they may be used to generate longer OBZCPs through certain recursive operations. For example, in 2003, Borwein and Ferguson carried out an exhaustive computer search to search new GCPs of lengths up to 100 [21]. Although some primitive OBZCPs are reported in [18], their lengths are capped to 25 only.
The main objective of this paper is thus to look for primitive optimal (or sub-optimal) OBZCPs of lengths $27\leq N\leq 49$. By leveraging the High Performance Computing (HPC) facility [^3] https://www.essex.ac.uk/staff/it-services/hpc. at the University of Essex, many new OBZCPs are found by a fine-grained computer search algorithm involving mapping, grouping, and Gray code counting. Besides the OBZCP optimality criteria on maximum ZCZ width and minimum out-of-zone AACF sums, we also use the Pursley-Sarwate criterion [22, 23] for selection of the best binary odd-length pairs with lowest combined auto-correlation and cross-correlation.
This paper is organized as follows. In Section II, we give some preliminaries and important equations/facts on OBZCPs. In Section III, we introduce the general considerations and programming techniques of the proposed algorithm. In Section IV, new OBZCPs found by computer search are presented. A list of OBZCPs with smallest demerit factors are given in Section V. Finally, this paper is concluded in Section VI.
@@ II-A Binary Z-complementary pairs (ZCPs)
In this case, $\rho_{\mathbf{a}}(\tau)+\rho_{\mathbf{b}}(\tau)$ for $Z\leq\tau\leq N-1$, is called the out-of-zone aperiodic auto-correlation sum of $\mathbf{a}$ and $\mathbf{b}$ at time-shift $\tau$. When $Z=N$, a ZCP reduces to a Golay complementary pair (GCP) [2].
Fact 1
@@
[…]
@@
II-B Pursley-Sarwate Criterion [22, 23]
@@
Furthermore, let us define the Pursley-Sarwate criterion of binary pair $(\mathbf{a},\mathbf{b})$ as follows:
@@
According to Pursley and Sarwate [22], we have $\text{PSC}(\mathbf{a},\mathbf{b})\geq 1$. Katz and Moore pointed out in [23] that binary GCPs of lengths $2^{\alpha}10^{\beta}26^{\gamma}$ (where $\alpha,\beta,\gamma$ are non-negative integers) satisfy the Pursley-Sarwate lower bound with equality, i.e., $\text{PSC}(\mathbf{a},\mathbf{b})=1$.
In this paper, we also use Pursley-Sarwate criterion as a sieve to select OBZCPs. Specifically, we look for optimal OBZCPs or $Z$-optimal OBZCPs with PSC values closest to 1.
@@ III-A General Considerations
In this section we describe our main considerations for our proposed algorithm with the complexity of $O(2^{N})$. Exhaustive computer search of all pairs is infeasible due to the high complexity of $O(2^{2N})$. We present a novel algorithm based on a two-step approach using the algebraic constructions and software data structures within a reasonable memory constraint.
@@
  • Negation: if $(\mathbf{a},\mathbf{b})$ is a solution, then so is $(neg(\mathbf{a}),\mathbf{b})$, where $neg(\mathbf{a})=(1-a_{0},1-a_{1},\ldots,1-a_{N-1})$;
  • Reciprocal: if $(\mathbf{a},\mathbf{b})$ is a solution, then so is $(recp(\mathbf{a}),\mathbf{b})$, where $recp(\mathbf{a})=(a_{N-1},\ldots,a_{1},a_{0})$.
Naturally, a series of these operations above lead to an equivalent solution. For instance $(neg(\mathbf{b})$, $recp(neg(\mathbf{a})))$ is also equivalent to $(\mathbf{a},\mathbf{b})$.
@@
$\left[\begin{matrix}a_{0}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{1}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{(N-3)/2}}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{a_{(N-1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{(N+1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{a_{N-2}}&1\\ a_{0}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{1}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{(N-3)/2}}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}{b_{(N-1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{(N+1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}{b_{N-2}}&1\end{matrix}\right].$
Next, we leverage the following three important data handling strategies in our proposed algorithm:
@@ III-B Mapping
The purpose of mapping is to link each sequence $\mathbf{b}$ with a proper sequence $\mathbf{a}$ such that the ZCZ can be achieved. Based on equation (3), we have
@@
for each $\tau=1,\ldots,(N-3)/2$. Suppose that there are $k$ sequences of $\mathbf{a}$ which satisfy the above equation. We store all the $\rho$ values and their corresponding sequences $\mathbf{a}$ in a map $\{\rho_{\mathbf{a}}\rightarrow[\mathbf{a}_{1},\ldots,\mathbf{a}_{k}]\}$ where
@@
for all $\mathbf{a}\in\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$ that generate the same $\rho$.
@@ III-C Grouping
Let us define a chunk
$\mathbf{C}=(c_{1},c_{2},...,c_{(N-3)/2)})$ (10)
and we say that $\mathbf{a}\in\mathbf{C}$ iff
$c_{i}=a_{i}\oplus a_{N-1-i},i=1,\ldots,(N-3)/2$ (11)
Thus, each $\mathbf{a}$ belongs to exactly one chunk. By noting that that $a_{0},a_{(N-1)/2},a_{N-1}$ are fixed, each chunk contains $2^{(N-3)/2}$ sequences of $\mathbf{a}$ and one needs to iterate over $2^{(N-3)/2}$ chunks.
@@ III-D Gray code counting
For every sequence $\mathbf{b}$, note that $b_{N-1}=1$, $b_{(N-1)/2}$ is fixed and $b_{0}=a_{0}$ (4). To proceed, $(b_{1},b_{2},\ldots,b_{(N-3)/2})$ is said to be the lower part of sequence $\mathbf{b}$ and $(b_{(N+1)/2},b_{(N+3)/2},\ldots,b_{(N-2)})$ its upper part. Having $2^{(N-3)/2}$ chunks, we need to iterate over the lower parts of the $2^{(N-3)/2}$ $\mathbf{b}$ sequences. Our key idea is to use a Gray code to represent the lower part, whereby the upper part is recalculated based on (4). A Gray code is characterized by having the next codeword with only one bit changed. This gives us two advantages:
@@
  • For such a change, the weight of sequence $\mathbf{b}$ is either increased or decreased by $2$ or stays unchanged. Hence, we can track the weight $q$ of sequence $\mathbf{b}$ without recalculating it for each modification. Having directly all possible $(p,q)$ pairs [18], we can directly exclude all the non-matching combinations.

### v3

# A Computer Search of New OBZCPs of Lengths up to 49

@@ Abstract
This paper aims to search for new optimal and sub-optimal Odd Binary Z-Complimentary Pairs (OBZCPs) for lengths up to 49. As an alternative to the celebrated binary Golay complementary pairs, optimal OBZCPs are the best almost-complementary sequence pairs having odd lengths. We introduce a computer search algorithm with time complexity $O(2^{N})$, where $N$ denotes the sequence length and then show optimal results for all $27\leq N\leq 33$ and $N=37,41,49$. For those sequence lengths (i.e., $N=35,39,43,45,47$) with no optimal pairs, we show OBZCPs with largest zero-correlation zone (ZCZ) widths (i.e., $Z$-optimal). Finally, based on the Pursley–Sarwate criterion (PSC), we present a table of OBZCPs with smallest combined auto-correlation and cross-correlation.
@@ I-A Background
Complementary pairs of sequences are useful in coding theory, wireless communication, radar sensing, and signal processing. The general design objective is to find two equal-length sequences whose maximum aperiodic auto-correlation function (AACF) sums are as small as possible (ideally zero). Pioneered by Marcel J. E. Golay [1] in 1951, binary complementary pairs were first studied in his design of infrared multislit spectrometry, a detector which isolates the desired radiation with a fixed single wavelength from background radiation with many different wavelengths. Formally, a pair of sequences is called a Golay complementary pair (GCP) [2] if their AACF sums are zero for all the non-zero time-shifts. Since it is generally difficult to find a single unimodular sequence with zero AACF sidelobes [^1] Although a Huffman sequence has zero AACF sidelobes except for the end time-shift, its sequence elements may have different magnitudes., GCP provides a solution by allowing two sequences to work in a collaborative way. Some representative applications of the GCPs (and their extensions/variants) include: peak-to-mean envelope power ratio (PMEPR) reduction of multicarrier signals [3, 4, 5, 6], Doppler resilient radar waveforms [7, 8, 9], channel estimation [10, 11, 12, 13], inter-cell interference rejection [14], multicarrier code-division multiple access [15, 16], etc.
Despite their wide applications, however, binary GCPs are limited to even-lengths only. More specifically, the existing binary GCPs are only known to have sequence lengths of $2^{\alpha}10^{\beta}26^{\gamma}$ (where $\alpha,\beta,\gamma$ are non-negative integers). For the odd-length case, optimal binary almost–complementary pairs are studied in [17]. Their optimization criteria is to maximize the zero-correlation zone (ZCZ) width of an odd-length binary pair, whilst minimizing its out-of-zone AACF sums. Such optimal binary pairs exhibit the correlation properties closest to GCPs and are called optimal odd-length binary Z-complementary pairs (OBZCPs). It is found that each optimal OBZCP has maximum ZCZ width of $(N+1)/2$, and minimum out-of-zone AACF sum magnitude of 2, where $N$ denotes the sequence length (odd) [17].
@@ I-B Contributions
Searching for optimal OBZCPs is of interest for understanding their deeper structural properties and for providing higher level of flexibility in practical applications. Besides the aforementioned systematic constructions, it is equally important to look for short OBZCPs because they may be used to generate longer OBZCPs through certain recursive operations. For example, in 2003, Borwein and Ferguson carried out an exhaustive computer search to search new GCPs of lengths up to 100 [21]. Although some OBZCPs are reported in [17], their lengths are capped to 25 only.
The main objective of this paper is thus to look for new optimal (or sub-optimal) OBZCPs of lengths $27\leq N\leq 49$. By leveraging the High Performance Computing (HPC) facility [^3] https://www.essex.ac.uk/staff/it-services/hpc. at the University of Essex, a number of new OBZCPs are found by a fine-grained computer search algorithm involving mapping, grouping, and Gray code counting. Besides the OBZCP optimality criteria on maximum ZCZ width and minimum out-of-zone AACF sums, we also use the Pursley–Sarwate criterion [22, 23] for selection of the best binary odd-length pairs with lowest combined auto-correlation and cross-correlation. For reproducibility of these results, our source code can be found at https://github.com/peterkazakov/obzcp.
This paper is organized as follows. In Section II, we give some preliminaries and important equations/facts on OBZCPs. In Section III, we introduce the general considerations and programming techniques of the proposed algorithm. In Section IV, new OBZCPs found by computer search are presented. A list of OBZCPs with smallest smallest combined auto-correlation and cross-correlation are given in Section V. Finally, this paper is concluded in Section VI.
@@ II-A Binary Z-complementary pairs (ZCPs)
In this case, $\rho_{\mathbf{a}}(\tau)+\rho_{\mathbf{b}}(\tau)$ for $Z\leq\tau\leq N-1$, is called the out-of-zone aperiodic auto-correlation sum of $\mathbf{a}$ and $\mathbf{b}$ at time-shift $\tau$. When $Z=N$, a ZCP reduces to a Golay complementary pair (GCP) [2]. An OBZCP refers to a binary ZCP with odd-length.
Fact 1.
@@
$\left\{\begin{array}[]{cl}a_{0}+a_{N-1}+b_{0}+b_{N-1}&\equiv~{}0~{}(\text{mod}~{}2),\\ a_{r}+a_{N-1-r}+b_{r}+b_{N-1-r}&\equiv~{}1~{}(\text{mod}~{}2),\end{array}\right.$ (4)
@@
II-B Pursley–Sarwate Criterion [22, 23]
@@
Furthermore, let us define the Pursley–Sarwate criterion of binary pair $(\mathbf{a},\mathbf{b})$ as follows:
@@
According to Pursley and Sarwate [22], we have $\text{PSC}(\mathbf{a},\mathbf{b})\geq 1$. Katz and Moore pointed out in [23] that binary GCPs of lengths $2^{\alpha}10^{\beta}26^{\gamma}$ (where $\alpha,\beta,\gamma$ are non-negative integers) satisfy the Pursley–Sarwate lower bound with equality, i.e., $\text{PSC}(\mathbf{a},\mathbf{b})=1$.
In this paper, we also use Pursley–Sarwate criterion as a sieve to select OBZCPs. Specifically, we look for optimal OBZCPs or $Z$-optimal OBZCPs with PSC values closest to 1.
@@ III-A General Considerations
In this section we describe our main considerations for our proposed algorithm with the time complexity of $O(2^{N})$. Exhaustive computer search of all pairs is infeasible due to the high complexity of $O(2^{2N})$. We present a novel algorithm based on a two-step approach using the algebraic constructions and software data structures within a reasonable memory constraint.
@@
  • Negation: if $(\mathbf{a},\mathbf{b})$ is a solution, then so is $(\operatorname{neg}(\mathbf{a}),\mathbf{b})$, where $\operatorname{neg}(\mathbf{a})=(1-a_{0},1-a_{1},\ldots,1-a_{N-1})$;
  • Reversal: if $(\mathbf{a},\mathbf{b})$ is a solution, then so is $(\operatorname{rvr}(\mathbf{a}),\mathbf{b})$, where $\operatorname{rvr}(\mathbf{a})=(a_{N-1},\ldots,a_{1},a_{0})$.
Naturally, a series of these operations above lead to an equivalent solution. For instance $(\operatorname{neg}(\mathbf{b})$, $\operatorname{rvr}(\operatorname{neg}(\mathbf{a})))$ is also equivalent to $(\mathbf{a},\mathbf{b})$.
@@
$\left[\begin{matrix}a_{0}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{a_{1}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{a_{(N-3)/2}}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{a_{(N-1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{a_{(N+1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{a_{N-2}}&1\\ a_{0}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{b_{1}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{b_{(N-3)/2}}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{b_{(N-1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{b_{(N+1)/2}}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\ldots}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{b_{N-2}}&1\end{matrix}\right].$
Next, we leverage the following three important data processing strategies in our proposed algorithm:
@@ III-B Mapping
The purpose of mapping is to link each sequence $\mathbf{b}$ with a proper sequence $\mathbf{a}$ such that the ZCZ can be achieved. Based on (3), we have
@@
for each $\tau=1,2,\ldots,(N-3)/2$. Suppose that there are $k$ sequences of $\mathbf{a}$ which satisfy the above equation. We store all the $\rho$ values and their corresponding sequences $\mathbf{a}$ in a map $\{\rho_{\mathbf{a}}\rightarrow[\mathbf{a}_{1},\ldots,\mathbf{a}_{k}]\}$ where
@@
for all $\mathbf{a}\in\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$ that generate the same $\rho$. We will denote by $\text{map}[\rho_{\mathbf{a}}]$ for the whole list of $\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$.
Example 1.
For $N=11$, $\rho_{\mathbf{a}}=(2,-1,2,1)$, we have
[…]
$~{}(1,0,1,1,0,0,1,1,1,1,1),$
$~{}(1,1,0,0,1,0,0,0,0,0,1),$
[…]
@@ III-C Grouping
For ease of presentation, denote by $N_{\mathbf{a}}$ the integer associated to the binary sequence $\mathbf{a}$ and vice versa.
For any binary vector $\mathbf{c}=(c_{1},c_{2},\dotsc,c_{(N-3)/2)})$, define a set $\mathbf{C}$ comprising of all the possible length-$N$ binary sequences $\mathbf{a}$, each satisfying
$c_{i}=a_{i}\oplus a_{N-1-i},i=1,2,\ldots,(N-3)/2$ (10)
where $\oplus$ is the binary xor operation. By noting that ${a}_{0},{a}_{(N-1)/2},{a}_{N-1}$ are fixed, each set $\mathbf{C}$ contains $2^{(N-3)/2}$ sequences of $\mathbf{a}$ and we have $2^{(N-3)/2}$ sets for all possible combinations of $\mathbf{c}$.
Example 2.
Consider $N=7$, $Z=(N+1)/2=4$, $\mathbf{c}=(1,0)$. One can generate the following $\mathbf{C}{(\mathbf{c})}$:
$\{(a_{0},0,0,a_{3},0,1,1),(a_{0},1,0,a_{3},0,0,1),$
$~{}(a_{0},0,1,a_{3},1,1,1),(a_{0},1,1,a_{3},1,0,1)\}.$
@@ III-D Gray code counting
For every sequence $\mathbf{b}$, note that $b_{N-1}=1$, $b_{(N-1)/2}$ is fixed and $b_{0}=a_{0}$ as shown in (4). To proceed, $(b_{1},b_{2},\ldots,b_{(N-3)/2})$ is said to be the lower part of sequence $\mathbf{b}$ and $(b_{(N+1)/2},b_{(N+3)/2},\ldots,b_{(N-2)})$ its upper part. Our key idea is to use a Gray code to represent the lower part, whereby the upper part is recalculated based on (4).
In the sequel, “codeword” and “sequence” may be used interchangeably. The weight of each codeword is equal to its number of 1’s. Comparison between codewords (sequences), e.g., $\mathbf{a}\geq\mathbf{b}$, is performed by comparing their associated integers, i.e., $N_{\mathbf{a}}\geq N_{\mathbf{b}}$. A Gray code is characterized by having the next codeword with only one bit changed. This gives us two advantages:
@@
  • For such a change, the weight of sequence $\mathbf{b}$ is either increased or decreased by $2$ or stays unchanged. Hence, we can track the weight of sequence $\mathbf{b}$ without recalculating it for each modification. By pre-calculating all the admissible weight pairs associated to $(\mathbf{a},\mathbf{b})$ (see Property 3 of [17]), we can directly exclude all the non-matching combinations.